OpenStudy (anonymous):

correct ? Simplify. (4x2 + 7x - 5) - (3x2 - 2x - 4 x^2 + 9x - 1

OpenStudy (experimentx):

correct

Directrix (directrix):

(4x2 + 7x - 5) - (3x2 - 2x - 4) = 4x^2 + 7x - 5 - 3x^2 + 2x + 4 = x^2 + 9x -1. Yes, your answer is correct.